# New Applications of Ergodic Theory to Sets of Differences

### v1

@@ 3.1 Ergodic sequences
Now, with the help of Theorem 3.1, one can use Theorem 3.1 to produce plenty of examples of norm ergodic sequences. For example, suppose that $f(x)\in\mathbb{R}[x]$ is a polynomial with two rationally independent coefficients. Then $\lambda f$ satisfies condition (ii) of Theorem 3.5 for all nonzero $\lambda\in\mathbb{R}$, thus $(\lambda f(n))_{n=1}^{\infty}$ is uniformly distributed mod 1, therefore $([f(n)])_{n=1}^{\infty}$ is norm ergodic by Theorem 3.4.
Lemma 3.7 in [10] is another result which affords us an assortment of norm ergodic sequences when combined with Theorem 3.1. Specifically, one finds that if $g\in\mathbf{U}$ and $k$ a positive integer such that $g(x)/x^{k}\rightarrow 0$ but $g(x)/x^{k+1}\rightarrow\infty$ as $x\rightarrow\infty$, then $([g(n)])_{n=1}^{\infty}$ is norm ergodic.
Some examples of norm ergodic sequences are:
$[\sqrt{2}n^{2}+\sqrt{3}n],\quad[n^{c}]\text{ with }c>0,c\notin\mathbb{N},\quad[n\log(n)],\quad\left[\frac{n^{2}}{\log(n)}\right],\quad[\log^{t}(n)]\text{ with }t>1$
for $n=1,2,3,\ldots$
It was shown in [8, Theorem B] that a rather large subclass of norm ergodic sequences guaranteed by [10] and Theorem 3.1 are also pointwise good. For example, the following sequence are pointwise good:
$n^{3}+n^{2}+n+1,\quad\left[n\log(n)\right],\quad\left[n^{c}\right]\text{ with }c>0,\quad\left[\frac{n^{2}}{\log(n)}\right]$
for $n=1,2,3,\ldots$
On the other hand, one can show that, while $([\log^{t}(n)])_{n=1}^{\infty}$ is norm ergodic for any $t>1$, it is not pointwise good As was mentioned above, if a sequence $(a_{n})_{n=1}^{\infty}$ in $\mathbb{N}$ is pointwise good and norm ergodic then it is pointwise ergodic. On the base of this, we can obtain many sequences which are pointwise ergodic. For example, the following sequences are pointwise ergodic:
$[\sqrt{2}n^{2}+\sqrt{3}n],\quad[n^{c}]\text{ with }c>0,c\notin\mathbb{N},\quad[n\log(n)],\quad\left[\frac{n^{2}}{\log(n)}\right]$ (3.9)
for $n=1,2,3,\ldots$
3.2 Proof of a General Variant of Theorem 1
The goal of this subsection is to present a proof of a general variant of Theorem 1, namely Theorem 3.2. The proof of Theorem 3.2 will rely on two theorems that we will presently formulate and prove.
Theorem 3.6 (cf. [7, Corollary 7.2]). Suppose that $(a_{n})_{n=1}^{\infty}$ is a norm ergodic sequence in $\mathbb{N}$. If $(X,\mathcal{B},\mu,T)$ is an invertible measure-preserving system and $A\in\mathcal{B}$ with $\mu(A)>0$, then
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{n=1}^{N}\mu(A\cap T^{a_{n}}A)\geq(\mu(A))^{2}.$ (3.10)
Proof.
By using weak convergence, if $T$ is ergodic we get that
@@ 3.2 Proof of a General Variant of Theorem 1
where $f:=1_{A}$. If $T$ is not ergodic then the inequality 3.10 follows by utilizing the ergodic decomposition of $\mu$ (for example see [6, Section 5]). ∎
@@
Proof.
Sets of the form $T^{-n}B$ with $n\in\mathbb{Z}$ form a countable collection of measurable subsets of $X$, so after possibly subtracting a set of measure zero from $X$ we may assume, without loss of generality, that if $n_{1},\ldots,n_{j}\in\mathbb{Z}$ such that $T^{-n_{1}}B\cap\cdots\cap T^{-n_{j}}B\neq\emptyset$, then $\mu\left(T^{-n_{1}}B\cap\cdots\cap T^{-n_{j}}B\right)>0$.
@@
$\lim_{N\rightarrow\infty}\int 1_{B}\cdot f_{N}=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{n=1}^{N}\mu(B\cap T^{-[g(n)]}B)\geq(\mu(B))^{2}.$ (3.13)
@@
$\int_{B}f=\int 1_{B}\cdot f=\int\limsup_{N\rightarrow\infty}(1_{B}\cdot f_{N})\geq\limsup_{N\rightarrow\infty}\int 1_{B}\cdot f_{N}=\lim_{N\rightarrow\infty}\int(1_{B}\cdot f_{N})\geq(\mu(B))^{2},$ (3.14)
@@
by equation 3.18. Now suppose that $n_{1},\ldots,n_{m}\in P$. Then for $i=1,\ldots,m$ we have $x_{0}\in B\cap T^{-[g(n_{i})]}B$, so $B\cap T^{-[g(n_{1})]}B\cap\cdots\cap T^{-[g(n_{m})]}B\neq\emptyset$, thus $\mu(B\cap T^{-[g(n_{1})]}B\cap\cdots\cap T^{-[g(n_{m})]}B)>0$. ∎
@@
Proof.
By Theorem 2, for each $i=1,\ldots,k$ there exists an invertible measure-preserving system $(X_{i},\mathcal{B}_{i},\mu_{i},T_{i})$ and a set $B_{i}\in\mathcal{B}_{i}$ with $\mu_{i}(B_{i})=\overline{d}(S_{i})$ such that $S_{i}$, $B_{i}$, and $T_{i}$ satisfy equation (2.2). Let $(X,\mathcal{B},\mu,T)$ be the product of the systems $(X_{i},\mathcal{B}_{i},\mu_{i},T_{i})$ for $1\leq i\leq k$, where $B=B_{1}\times\dots\times B_{k}$ (see the proof of Theorem 1 for an explicit description of $(X,\mathcal{B},\mu,T)$). By Theorem 3.2 there exists a set $S\subseteq\mathbb{N}$ with $d(S)\geq\mu(B)$ such that for any $n_{1},\ldots,n_{r}\in S$ we have $\mu(B\cap T^{-[g(n_{1})]}B\cap\cdots\cap T^{-[g(n_{r})]}B)>0$. Then for $1\leq i\leq k$ and for all $n_{1},\dots,n_{r}\in S$ we have
@@
$\geq\prod_{i=1}^{k}\mu_{i}(B_{i}\cap(B_{i}-[g(n_{1})])\cap\dots\cap(B_{i}-[g(n_{r})]))$
$=\mu(B_{i}\cap(B_{i}-[g(n_{1})])\cap\dots\cap(B_{i}-[g(n_{r})]))>0.$
@@
$\Delta_{1,g}(S)=\{n\in\mathbb{N}:S\cap(S-[g(n)])\neq\emptyset\},$
$\Delta_{2,g}(S)=\{n\in\mathbb{N}:\overline{d}(S\cap(S-[g(n)]))>0\}.$
Theorem 3.10. Let $g:\mathbb{N}\rightarrow[1,\infty)$ and assume $([g(n)])_{n=1}^{\infty}$ is pointwise ergodic. If $S_{1},\dots,S_{k}\subseteq\mathbb{N}$ satisfy $\overline{d}(S_{i})>0$ for each $i=1,\dots,k$, then there is $S\subseteq\mathbb{N}$ with $d(S)\geq\prod_{i=1}^{k}\overline{d}(S_{i})$ and $\Delta_{1,g}(S)\subseteq D_{g}:=\bigcap_{i=1}^{k}\Delta_{2,g}(S_{i})$. Furthermore, we have that $\overline{d}(D_{g})\geq\prod_{i=1}^{k}\overline{d}(S_{i})$.
Proof.
By Theorem 1 there exists $S\subseteq\mathbb{N}$ with $d(S)\geq\prod_{i=1}^{k}\overline{d}(S_{i})$ and $\Delta_{1}(S)\subseteq\bigcap_{i=1}^{k}\Delta_{2}(S_{i})$. If $n\in\Delta_{1,g}(S)$ then $[g(n)]\in\Delta_{1}(S)$, hence $[g(n)]\in\bigcap_{i=1}^{k}\Delta_{2}(S_{i})$ since $\Delta_{1}(S)\subseteq\bigcap_{i=1}^{k}\Delta_{2}(S_{i})$, therefore $n\in\bigcap_{i=1}^{k}\Delta_{2,g}(S_{i})$. This shows that $\Delta_{1,g}(S)\subseteq D_{g}$.
Next, by Corollary 3.2, there is a set $P\subseteq\mathbb{N}$ with $d(P)\geq\prod_{i=1}^{k}\overline{d}(S_{i})$ such that, in particular, for all $n\in P$ and for each $1\leq i\leq k$ we have $\overline{d}(S_{i}\cap(S_{i}-[g(n)]))>0$. It therefore follows that $P\subseteq D_{g}$, hence $\overline{d}(D_{g})\geq d(P)\geq\prod_{i=1}^{k}\overline{d}(S_{i})$. ∎

### v2

@@ 3.1 Ergodic sequences
Now, with the help of Theorem 3.1, one can use Theorem 3.1 to produce plenty of examples of norm ergodic sequences. For example, $[\log^{t}(n)]\text{ with }t>1$ is norm ergodic. Also, if $g\in\mathbf{U}$ and $k\in\mathbb{N}$ are such that $g(x)/x^{k}\rightarrow\infty$ but $g(x)/x^{k+1}\rightarrow 0$ as $x\rightarrow\infty$, then $\lambda g$ satisfies condition (ii) of Theorem 3.1 for all nonzero $\lambda\in\mathbb{R}$ (this follows from repeated applications of L’Hospital), and so $([g(n)])_{n=1}^{\infty}$ is norm ergodic by Theorem 3.1. Some examples obtained in this way are
$[n^{c}]\text{ with }c>0,c\notin\mathbb{N},\quad[n\log(n)],\quad\left[\frac{n^{2}}{\log(n)}\right],\quad\quad\text{for }n=1,2,3,\ldots$
To describe another class of examples, suppose that $f(x)\in\mathbb{R}[x]$ is a polynomial with two coefficients other than the constant term, which are linearly independent over $\mathbb{Q}$. Then $\lambda f$ satisfies condition (ii) of Theorem 3.1 for all nonzero $\lambda\in\mathbb{R}$, and so $([f(n)])_{n=1}^{\infty}$ is norm ergodic by Theorem 3.1. So, for instance, the sequence $([\sqrt{2}n^{2}+\sqrt{3}n])_{n=1}^{\infty}$ is norm ergodic.
It was shown in [12, Theorem B] that many sequences guaranteed to be norm ergodic by [14] and Theorem 3.1 are also pointwise good. For example, the following sequences are pointwise good:
$n^{3}+n^{2}+n+1,\quad\left[n\log(n)\right],\quad\left[n^{c}\right]\text{ with }c>0,\quad\left[\frac{n^{2}}{\log(n)}\right],\quad\quad\text{for }n=1,2,3,\ldots$
On the other hand, one can show that, while $([\log^{t}(n)])_{n=1}^{\infty}$ is norm ergodic for any $t>1$, it is not pointwise good (this follows from [20, Theorem 2.16]; see [20, Example 2.18] for a proof of the case $t=2$).
As was mentioned above, if a sequence $(a_{n})_{n=1}^{\infty}$ in $\mathbb{N}$ is pointwise good and norm ergodic, then it is pointwise ergodic. On the base of this, we can obtain many sequences which are pointwise ergodic. For example, the following sequences are pointwise ergodic:
$[\sqrt{2}n^{2}+\sqrt{3}n],\quad[n^{c}]\text{ with }c>0,c\notin\mathbb{N},\quad[n\log(n)],\quad\left[\frac{n^{2}}{\log(n)}\right],\quad\quad\text{for }n=1,2,3,\ldots$ (3.9)
3.2 Proofs of Theorems 3.2 and 3.2
The goal of this subsection is to present a proof of a general variant of Theorem 1, namely Theorem 3.2. The proof of Theorem 3.2 will rely on two results that we will presently formulate and prove.
Theorem 3.6 (cf. [9, Corollary 7.2]). Suppose that $(a_{n})_{n=1}^{\infty}$ is a norm ergodic sequence in $\mathbb{N}$. If $(X,\mathcal{B},\mu,T)$ is an invertible measure-preserving system and $A\in\mathcal{B}$ with $\mu(A)>0$, then
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{n=1}^{N}\mu(A\cap T^{a_{n}}A)\geq\mu(A)^{2}.$ (3.10)
Proof. By using weak convergence, if $T$ is ergodic, it follows from the assumption that $(a_{n})$ is a norm ergodic sequence that
@@
If $T$ is not ergodic, then the inequality 3.10 follows by utilizing the ergodic decomposition of $\mu$ (an example of such an application of the ergodic decomposition is delineated in [8, Section 5, page 48]). $\square$
@@
Proof. We will employ an argument similar to the one used in the proof of [5, Theorem 1.2], cited above as Theorem 2.
After deleting (if needed) a set of measure zero from each $T^{-n}B$, we can and will assume without loss of generality that for any $n_{1},...,n_{j}\in\mathbb{Z}$, if $T^{-n_{1}}B\cap\cdots\cap T^{-n_{j}}B\neq\emptyset$, then $\mu(T^{-n_{1}}B\cap\cdots\cap T^{-n_{j}}B)>0$.
@@
$\lim_{N\rightarrow\infty}\int 1_{B}\cdot f_{N}=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{n=1}^{N}\mu(B\cap T^{-[g(n)]}B)\geq\mu(B)^{2}.$ (3.13)
@@
$\int_{B}f=\int 1_{B}\cdot f=\int\limsup_{N\rightarrow\infty}(1_{B}\cdot f_{N})\geq\limsup_{N\rightarrow\infty}\int 1_{B}\cdot f_{N}=\lim_{N\rightarrow\infty}\int(1_{B}\cdot f_{N})\geq\mu(B)^{2},$ (3.14)
@@
by formula 3.18. Now suppose that $n_{1},\ldots,n_{m}\in P$. Then for $i=1,\ldots,m$ we have $x_{0}\in B\cap T^{-[g(n_{i})]}B$, so $B\cap T^{-[g(n_{1})]}B\cap\cdots\cap T^{-[g(n_{m})]}B\neq\emptyset$, thus $\mu(B\cap T^{-[g(n_{1})]}B\cap\cdots\cap T^{-[g(n_{m})]}B)>0$. $\square$
@@
Proof. By Theorem 2, for each $i=1,\ldots,k$ there exists an invertible measure-preserving system $(X_{i},\mathcal{B}_{i},\mu_{i},T_{i})$ and a set $B_{i}\in\mathcal{B}_{i}$ with $\mu_{i}(B_{i})=\overline{d}(S_{i})$ such that $S_{i}$, $B_{i}$, and $T_{i}$ satisfy formula (2.2). Let $(X,\mathcal{B},\mu,T)$ be the product of the systems $(X_{i},\mathcal{B}_{i},\mu_{i},T_{i})$ for $1\leq i\leq k$, where $B=B_{1}\times\dots\times B_{k}$ (see the proof of Theorem 1 for an explicit description of $(X,\mathcal{B},\mu,T)$). By Theorem 3.2 there exists a set $S\subseteq\mathbb{N}$ with $d(S)\geq\mu(B)$ such that for any $n_{1},\ldots,n_{r}\in S$ we have $\mu(B\cap T^{-[g(n_{1})]}B\cap\cdots\cap T^{-[g(n_{r})]}B)>0$. Then for any $1\leq i\leq k$ and for all $n_{1},\dots,n_{r}\in S$, we have
@@
$\geq\prod_{i=1}^{k}\mu_{i}(B_{i}\cap T_{i}^{-[g(n_{1})]}B_{i}\cap\dots\cap T_{i}^{-[g(n_{r})]}B_{i})=\mu(B\cap T^{-[g(n_{1})]}B\cap\dots\cap T^{-[g(n_{r})]}B)>0.$
$\square$
@@
$\Delta_{1,g}(S)=\{n\in\mathbb{N}:S\cap(S-[g(n)])\neq\emptyset\},\quad\Delta_{2,g}(S)=\{n\in\mathbb{N}:\overline{d}(S\cap(S-[g(n)]))>0\}.$
Theorem 3.10. Let $g:\mathbb{N}\rightarrow[1,\infty)$ and assume $([g(n)])_{n=1}^{\infty}$ is pointwise ergodic. If $S_{1},\dots,S_{k}\subseteq\mathbb{N}$ satisfy $\overline{d}(S_{i})>0$ for each $i=1,\dots,k$, then there is $S\subseteq\mathbb{N}$ with $d(S)\geq\prod_{i=1}^{k}\overline{d}(S_{i})$ and $\Delta_{1,g}(S)\subseteq D_{g}:=\bigcap_{i=1}^{k}\Delta_{2,g}(S_{i})$. Furthermore, we have $\overline{d}(\Delta_{1,g}(S))\geq\prod_{i=1}^{k}\overline{d}(S_{i})$. If, in addition, the function $g$ comes from a Hardy field with the property that for some $\ell\in\mathbb{N}\cup\{0\}$, $\lim_{x\to\infty}g^{(\ell)}(x)=\pm\infty$ and $\lim_{x\to\infty}g^{(\ell+1)}(x)=0$, then the set $D_{g}$ is thick.
Proof. By Theorem 1 there exists $S\subseteq\mathbb{N}$ with $d(S)\geq\prod_{i=1}^{k}\overline{d}(S_{i})$ and $\Delta_{1}(S)\subseteq\bigcap_{i=1}^{k}\Delta_{2}(S_{i})$. If $n\in\Delta_{1,g}(S)$, then $[g(n)]\in\Delta_{1}(S)\subseteq\bigcap_{i=1}^{k}\Delta_{2}(S_{i})$,which implies $n\in\bigcap_{i=1}^{k}\Delta_{2,g}(S_{i})$. This shows that $\Delta_{1,g}(S)\subseteq D_{g}$.
Next, by Corollary 3.2, there is a set $P\subseteq\mathbb{N}$ with $d(P)\geq d(S)\geq\prod_{i=1}^{k}\overline{d}(S_{i})$ such that, for all $n\in P$ we have $\overline{d}(S\cap(S-[g(n)]))>0$. It follows that $P\subseteq\Delta_{1,g}(S)$, hence $\overline{d}(\Delta_{1,g}(S))\geq d(P)\geq\prod_{i=1}^{k}\overline{d}(S_{i})$.
Now assume in addition that the function $g$ comes from a Hardy field with the property that for some $\ell\in\mathbb{N}\cup\{0\}$, $\lim_{x\to\infty}g^{(\ell)}(x)=\pm\infty$ and $\lim_{x\to\infty}g^{(\ell+1)}(x)=0$. Then [13, Theorem A] implies, via Furstenberg’s correspondence principle (Theorem 2), that the set $D_{g}$ is thick. $\square$
@@
We conclude this subsection by formulating a variant of Theorem 3.2 in which the condition of pointwise ergodicity of $([g(n)])$ is relaxed, which in turn extends the range of applicability of the result. But first, we need to introduce yet another definition.
Definition 3.11. Let $(a_{n})_{n=1}^{\infty}$ be a sequence in $\mathbb{N}$.
  • i)
    $(a_{n})$ is an averaging sequence of recurrence if for every invertible measure-preserving system $(X,\mathcal{B},\mu,T)$ and any set $A\in\mathcal{B}$ with $\mu(A)>0$,
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{n=1}^{N}\mu(A\cap T^{a_{n}}A)>0.$ (3.20)
  • ii)
    $(a_{n})$ is a uniform averaging sequence of recurrence if for every invertible measure-preserving system $(X,\mathcal{B},\mu,T)$ and any set $A\in\mathcal{B}$ with $\mu(A)>0$,
    $\lim_{N-M\rightarrow\infty}\frac{1}{N-M}\sum_{n=M+1}^{N}\mu(A\cap T^{a_{n}}A)>0.$ (3.21)
Note that, by Theorem 3.2, any sequence which is norm ergodic is also an averaging sequence of recurrence. For example, $([n^{c}])_{n=1}^{\infty}$ for $c>0$ is an averaging sequence of recurrence. On the other hand, one can show that when $c>0$, $c\notin\mathbb{N}$, the sequence $([n^{c}])_{n=1}^{\infty}$ is not a uniform averaging sequence of recurrence. A good supply of examples of uniform averaging sequences of recurrence is provided by polynomials. For example, if $q(n)\in\mathbb{Q}[n]$ is such that $q(\mathbb{Z})\subseteq\mathbb{Z}$ and is intersective, that is, if $\{q(n):n\in\mathbb{Z}\}\cap k\mathbb{Z}$ is non-empty for each $k\in\mathbb{N}$, then $(q(n))_{n=1}^{\infty}$ is a uniform averaging sequence of recurrence (see [10, Theorem 1.2]). Also, if $p(n)\in\mathbb{R}[n]$ has at least two coefficients other than the constant term which are linearly independent over $\mathbb{Q}$, then $([p(n)])_{n=1}^{\infty}$ is a uniform averaging sequence of recurrence (see [10, Example 1.3]).
Theorem 3.12. Let $g:\mathbb{N}\rightarrow[1,\infty)$ and assume that $([g(n)])_{n=1}^{\infty}$ is an averaging sequence of recurrence and is pointwise good. If $S_{1},\dots,S_{k}\subseteq\mathbb{N}$ satisfy $\overline{d}(S_{i})>0$ for each $i=1,\dots,k$, then there is $S\subseteq\mathbb{N}$ with $d(S)>0$ and $\Delta_{1,g}(S)\subseteq D_{g}:=\bigcap_{i=1}^{k}\Delta_{2,g}(S_{i})$. Furthermore, we have that $\overline{d}(\Delta_{1,g}(S))\geq d(S)>0$. If, in addition, the sequence $([g(n)])_{n=1}^{\infty}$ is a uniform averaging sequence of recurrence, then the set $D_{g}$ is syndetic.
Proof. Admittedly, the proof of Theorem 3.2 has similarities to, and uses ideas from, the proofs of Theorem 3.2, Theorem 3.2, and Corollary 3.2. However, the proof is not just a “mechanical” composition of the aforementioned proofs. In particular, the proof below has novel elements when it comes to showing that under the assumptions of Theorem 3.2, the set $D_{g}$ is syndetic.
By Theorem 2, for every $i=1,...,k$, there exists an invertible measure-preserving system $(X_{i},\mathcal{B}_{i},\mu_{i},T_{i})$ and a set $B_{i}\in\mathcal{B}_{i}$ with $\mu_{i}(B_{i})=\overline{d}(S_{i})$ satisfying formula (2.2). Let $(X,\mathcal{B},\mu,T)$ be the product of the systems $(X_{i},\mathcal{B}_{i},\mu_{i},T_{i})$, $i=1,...,k$ and let $B=B_{1}\times\cdots\times B_{k}$.
We will utilize a method similar to the one used in the proof of Theorem 3.2 to obtain the set $S$. Consider the family of sets $T^{-n}B$ with $n\in\mathbb{Z}$. After deleting (if needed) a set of measure zero from each $T^{-n}B$, we can and will assume without loss of generality that for any $n_{1},...,n_{j}\in\mathbb{Z}$, if $T^{-n_{1}}B\cap\cdots\cap T^{-n_{j}}B\neq\emptyset$, then $\mu(T^{-n_{1}}B\cap\cdots\cap T^{n_{j}}B)>0$. Now for all $N\in\mathbb{N}$, define $f_{N}=\frac{1}{N}\sum_{n=1}^{N}1_{T^{-[g(n)]}B}$ and put $f=\limsup_{N\to\infty}f_{N}$. Since $([g(n)])_{n=1}^{\infty}$ is an averaging sequence of recurrence, it follows that
$\lim_{N\to\infty}\int 1_{B}\cdot f_{N}=\lim_{N\to\infty}\frac{1}{N}\sum_{n=1}^{N}\int 1_{B}\cdot 1_{T^{-[g(N)]}B}=\lim_{N\to\infty}\frac{1}{N}\sum_{n=1}^{N}\mu(B\cap T^{-[g(n)]}B)>0.$ (3.22)
Since $1_{B}\cdot f=\limsup_{N\to\infty}(1_{B}\cdot f_{N})$, Fatou’s lemma implies
$\int_{B}f=\int 1_{B}\cdot f\geq\limsup_{N\to\infty}\int 1_{B}\cdot f_{N}=\lim_{N\to\infty}\int 1_{B}\cdot f_{N}>0.$ (3.23)
Thus, $\mu(B\cap\left\{f>0\right\})>0$. Since $([g(n)])_{n=1}^{\infty}$ is pointwise good, the limit
$\lim_{N\to\infty}f_{N}(x)=\lim_{N\to\infty}\frac{1}{N}\sum_{n=1}^{N}1_{T^{-[g(n)]}B}(x)$ (3.24)
exists for almost every $x\in X$. In particular, there exists $x_{0}\in B$ for which we have $f(x_{0})=\lim_{N\to\infty}f_{N}(x_{0})>0$. Now let $S=\left\{n\in\mathbb{N}:x_{0}\in B\cap T^{-[g(n)]}B\right\}$ and observe that (since $x_{0}\in B$)
$d(S)=\lim_{N\to\infty}\frac{\left|S\cap\left\{1,...,N\right\}\right|}{N}=\lim_{N\to\infty}\frac{1}{N}\sum_{n=1}^{N}1_{B\cap T^{-[g(n)]}B}(x_{0})=\lim_{N\rightarrow\infty}f_{N}(x_{0})=f(x_{0})>0.$ (3.25)
Moreover, for any $n_{1},...,n_{j}\in S$, we have $x_{0}\in B\cap T^{-[g(n_{1})]}B\cap\cdots\cap T^{-[g(n_{j})]}B$, which implies $\mu(B\cap T^{-[g(n_{1})]}B\cap\cdots\cap T^{-[g(n_{j})]}B)>0$ in light of our earlier assumptions. Therefore, by invoking Furstenberg’s Correspondence Principle (Theorem 2), we have that for any $\ell=1,...,k$,
$\overline{d}(S_{\ell}\cap(S_{\ell}-[g(n_{1})])\cap\cdots\cap(S_{\ell}-[g(n_{j})])\geq\prod_{i=1}^{k}\overline{d}(S_{i}\cap(S_{i}-[g(n_{1})])\cap\cdots\cap(S_{i}-[g(n_{j})])$
$\geq\prod_{i=1}^{k}\mu_{i}(B_{i}\cap T^{-[g(n_{1})]}B_{i}\cap\cdots\cap T^{-[g(n_{j})]}B_{i})=\mu(B\cap T^{-[g(n_{1})]}B\cap\cdots\cap T^{-[g(n_{j})]}B)>0.$
In particular, for any $n\in S$ and $\ell=1,...,k$, one has $\overline{d}(S_{\ell}\cap(S_{\ell}-[g(n)]))>0$. It follows that $\Delta_{1,g}(S)\subseteq D_{g}$.
Next, by Corollary 3.2, there is a set $P\subseteq\mathbb{N}$ with $d(P)\geq d(S)>0$ such that, for all $n\in P$ we have $\overline{d}(S\cap(S-[g(n)]))>0$. It follows that $P\subseteq\Delta_{1,g}(S)$, hence $\overline{d}(\Delta_{1,g}(S))\geq d(P)\geq d(S)>0$.
Lastly, suppose that $([g(n)])$ is a uniform averaging sequence of recurrence. By Furstenberg’s correspondence principle, there is an invertible measure-preserving system $(Y,\mathcal{A},\nu,U)$ and a set $A\in\mathcal{A}$ with $\nu(A)=d(S)$ such that, for all $n_{1},\ldots,n_{r}\in\mathbb{N}$,
$\overline{d}(S\cap(S-n_{1})\cap\cdots\cap(S-n_{r}))\geq\nu(A\cap U^{-n_{1}}A\cap\cdots\cap U^{-n_{r}}A).$
It follows that $1_{\Delta_{2,g}(S)}(n)\geq\overline{d}(S\cap(S-[g(n)]))\geq\nu(A\cap U^{-[g(n)]}A)$ for all $n\in\mathbb{N}$, and so
$\liminf_{N-M\rightarrow\infty}$ $\frac{|\Delta_{1,g}(S)\cap\{M+1,\ldots,N\}|}{N-M}=\liminf_{N-M\rightarrow\infty}\frac{1}{N-M}\sum_{n=M+1}^{N}1_{\Delta_{1,g}(S)}(n)$
$\geq\liminf_{N-M\rightarrow\infty}\frac{1}{N-M}\sum_{n=M+1}^{N}1_{\Delta_{2,g}(S)}(n)\geq\lim_{N-M\rightarrow\infty}\frac{1}{N-M}\sum_{n=M+1}^{N}\nu(A\cap U^{-[g(n)]}A)>0.$
This shows that $\Delta_{1,g}(S)$ is syndetic, hence $D_{g}$ is as well since $\Delta_{1,g}(S)\subseteq D_{g}$. $\square$
3.3 Some subtle points about sets of the form $D_{g}$ in Theorems 3.2 and 3.2